OpenStudy (anonymous):

What are the steps to solve this? (7.5-?)/2= -.85

OpenStudy (anonymous):

(7.5-x)/2=-.85 7.5-x=-.85(2) 7.5-x=-1.7 7.5+1.7=x 9.2=x